OpenStudy (anonymous):

What is the area of sector BEC when CE = 9 yd? HELPPP!

OpenStudy (campbell_st):

well I think you need a sector angle.. to be able to complete the question

OpenStudy (anonymous):

let me draw it

OpenStudy (anonymous):



OpenStudy (campbell_st):

ummm... you have a right angle and 120 in the same sector...?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

well I can understand why the question is difficult...  if you can find x then the sector area is \[A = \frac{x}{360} \times \pi \times 9^2\] hope it helps